OpenStudy (anonymous):

lim x->infinity of f(x) = x^2 - x^4

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

any way to show mathmatically the intuitive negative infinity?

OpenStudy (misty1212):

guess, i bet you will be right if it is not obvious, try \(x=100\) and see what you get

OpenStudy (anonymous):

right, so infinity minus infinity squared is negative infinity

OpenStudy (misty1212):

there is no such thing as infinity minus infinity squared

OpenStudy (anonymous):

so, how would you write that?

OpenStudy (misty1212):

not sure what you mean writing things like \(\infty-\infty\) is just a from, a shorthand for limits

OpenStudy (anonymous):

i guess I'm asking you to prove what is obvious via algebra

OpenStudy (misty1212):

oh i see

OpenStudy (misty1212):

yo would show (maybe with some algebra that given any \(N>0\) there is an \(x\) such that \[x^2-x^4<-N\]

OpenStudy (anonymous):

sounds like profs may just accept intuition in this case...

OpenStudy (michele_laino):

Please note that I can rewrite your function as below: \[x ^{2}-x ^{4}=x ^{4}\left( \frac{ 1 }{ x ^{2} }-1 \right)\]

OpenStudy (anonymous):

so, always a good idea to factor out highest degree of x?

OpenStudy (michele_laino):

yes! I think it is the standard procedure

OpenStudy (anonymous):

how do you get from there to negative infinity?

OpenStudy (anonymous):

oh, duh, I see it now, (0-1) * infinity

OpenStudy (michele_laino):

please note that x^4 is not an undetermined form, and also (1/x^2 -) is not an undetermined form, so their limit exist and you can apply the rule of multiplication of limits

OpenStudy (michele_laino):

opps...also (1/x^2 -1) is not an undetermined form

OpenStudy (anonymous):

thanks all for helping

OpenStudy (michele_laino):

thanks!